OpenStudy (anonymous):

Find the exact radian measure of the angle 120 degrees and 210 degrees

OpenStudy (anonymous):

remember that 180 is pi

OpenStudy (anonymous):

120/180 = 2/3 -> 2pi/3 210/180 = 7/6 -> 7pi/6

OpenStudy (anonymous):

so you have to know which fraction of pi it is

OpenStudy (anonymous):

120/180 = 2/3

OpenStudy (anonymous):

2/3 pi

OpenStudy (shadowfiend):

chaise -- reminders are useful to someone who isn't skilled.

OpenStudy (shadowfiend):

Let's not talk about that anymore. Solid answer imranmeah.

OpenStudy (anonymous):

thanks, shadow,